OpenStudy (anonymous):

A car is driving at a speed of 50 miles per hour. What is its speed in m/s?

OpenStudy (anonymous):

1 mile =1609meters \[\frac{(50miles)}{(1hr)}*\frac{(1609meters)}{(1 mile)}*\frac{(1 hr)}{(3600seconds)} = \text{22.35 meters per second}\]